Operations Report – December 12th

Crew 200 Operations Report 12-Dec-2018

SOL: 10

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5 pm last night to 10:40 am (SOC was still
at 78% this morning, not changing from yesterday night). On tonight at
5:30pm.

Solar— SOC 72%
Diesel Reading – 100 % (refilled)
Propane Reading – 30 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 350 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 3

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: yes (by media)
Hours: 60.0
Beginning charge: 100%
Ending charge: 54%
Currently charging: yes

Opportunity rover used: yes (by media)
Hours: 41.8
Beginning charge: 100%
Ending charge: 88%
Currently charging: yes

Curiosity rover used: yes (by media)
Hours: not checked
Beginning charge: 100%
Ending charge: not checked
Currently charging: yes

Notes on rovers: Rovers were used today by the media, we could not use
Opportunity for our planned EVA as they were using it. Curiosity was
not present at the base when we came back from EVA, where we checked
the batteries levels.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now
parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: This afternoon and evening are very windy,
we had to securize some of the tunnels tarps with zips. One on the RAM
tunnel had to be removed and placed in the RAM. Temperature in the hab
dropped from 23°C to 19°C, heater is working, it is improving.

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this morning to
study the radio coverage, everything was alright. Ventilation of
spacesuit 10 sometimes stops before starting later. Spacesuit 3 still
not functional.

Summary of Hab operations: we catched a mouse yesterday night, which
we released today far from the Hab during EVA.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal.
Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

EVA Report – December 12th

Crew 200 EVA Report 12-12-2018

EVA #5

Author: Antoine Bocquier

Purpose of EVA: 1) Map out locations that allow radio communication
between EVA crew and HabCom around Phobos Peak 2) Test how EVAs can
safely and effectively be conducted with three-member crew, including
rover use.

Start time: 10:35

End time: 12:30

Narrative: The media filming around the Hab this morning, we had to
delay a bit our planned EVA. In addition, in agreement with Shannon
they were using the rovers, so we had to change our plans and perform
a walking EVA. We chose to go to Phobos Peak, which we could not reach
on our EVA monday, to map the radio coverage.
The walking went very fine (30mn to get to Phobos Peak), once there,
we walked north up to Mountain Goat Trail before coming back to the
Hab by the north. Radio link was pretty clear all along, we estimated
that Phobos Peak could be a very good location to settle a radio relay
(to explore eastern areas) and may be accessible on its northern side.
We are currently processing the acquired data (radio check every 2mn +
GPS positions) to complete our radio coverage map.

Destination: Phobos Peak + Southern part of Mountain Goat Trail

Coordinates (use UTM NAD27 CONUS): 519500 E – 4250500 N

Participants: Lindsay, Antoine

Road(s) and routes per MDRS Map : N/A

Mode of travel: walking

Sol Summary – December 12th

Crew 200 Sol Summary 12Dec2018

Crew 200 Sol Summary Report 12-DEC-2018

Sol: 10

Summary Title: Exploring Phobos Peak and examining unusual objects
collected during EVA

Author’s name: Lindsay Rutter

Mission Status: All crew members are happy and healthy

Sol Activity Summary: Each morning, the first object I see when I look
out the east window is the silhouette of Phobos Peak Rock as the
rising sun lights it up from behind. We had a beautiful sunrise this
morning and Phobos Peak Rock emerged all the more majestic than usual.
I was happy to remember that it was our approved geographical
destination for our EVA. We conducted another 3 crew EVA and walked to
Phobos Peak. We then explored its surrounding areas, discovering a
curious aluminum object near the peak, which we returned to the
habitat for investigation. Winds became violent shortly after our
return, but we celebrated our successful EVA safely inside our habitat
with a dish of quinoa peppered with oregano from the GreenHab.

Look Ahead Plan: We will continue conducting 3 crew EVAs and mapping
radio strength and connectivity. We will also continue our
collaborative projects.

Anomalies in work: Heavy winds caused minor damage in the tunnels,
which we repaired. Heating system in the habitat has been working on
and off; it appears to be working now but we will continue monitoring
it.

Weather: Snow melting; clear in the morning but clouds and heavy winds
developed in the early afternoon.

Crew Physical Status: No crew members reported physical problems today.

EVA: A walking EVA was conducted at and around Phobos Peak. We
continued our communication protocol to assess radio strength and
connectivity in areas around the habitat. Our EVA was almost exactly
two hours.

Reports to be file: Operations Report, EVA Report, EVA Request,
GreenHab Report, Pictures Report, Journalist Report, Sol Report

Support Requested: N/A

Journalist Report – December 12th

As we prepare for breakfast, we decide on which Rovers we would like
to take on today’s EVA. I look out the window. The rovers are gone. I
suppose the robots are using them…

So change of plans. We have to do a walking EVA. Well, THEY do. It’s
my turn to do HabCom!I’ve been looking forward to this. Not because I
don’t like to walk around in a clunky spacesuit (I’d be on the wrong
planet if I did), but because I can get a different perspective as an
observer and director. I get Antoine and Lindsay set up with their
gloves and comms, squirming around their ECLSS backpacks. I allow them
to enter the airlock and begin depressurization. I look above through
the window to see them walk down the balcony and off to Phobos Peak.

It’s a rather boring HabCom. All of the places the team went to had at
least comprehend-able connection, though I suppose it is good to know
where we still have great coverage. I’m happy with my team, conducting
another successful mission.

The day is spreading by, the days are moving faster. Is tomorrow
Thursday? Is it really that close? I think of my home, my parents
eager to see me for Christmas. My friends who are done with finals and
can finally hang out. I feel the crisp chill of a Texas winter. It was
a distant thought only a few days ago, but now it’s nearly tangible.

I wonder if I will ever be here again. If this moment on the dusty
planet will be my last.
Even more pressing, I wonder if my work here will make a difference.

“You are clear to enter the airlock”

An abrupt wind storm sweeps the region. The tunnels that connect the
secure building are shaking. Antoine goes out to fix it. Realizing the
severity, he calls for help. Lindsay starts to put on her coat. Do I
go? Should I stay? What if the tunnel rips apart and I’m the only one
alive? I end up following them after a few minutes. The wind is strong
and rattling the whole structure. We make the decision to come back
inside, it’d be better to wait the storm than be so vulnerable.

Mars is not safe. There are people back home depending on our sense of
precaution and protocol to come back home. Moments like these make me
think of the real scenario, if I can make the calls to make sure that
happens.

Sometimes, I hope Mars will be boring.

Makiah Eustice

Greenhab Report – December 12th

Crew 200 Greenhab Report 12-DEC-2018
Greenhab Officer:
Makiah Eustice
Environmental Control:
Heating
40% Shade Cloth on
Average Temperatures:
Low: 24 C
High: 54 C
Hours of Supplemental Light:
5
Daily Water Usage of Crops:
8 gallons
Water in Blue Tank (lbs):
68%
Times of Watering for Crops:
1500
Changes to crops:
NONE
Narrative:
Everything seems to be operating normally.
Harvest:
1.1 g Oregano
Support/Supplies Needed:
N/A